# label=0.,leftmargin=15pt,labelwidth=10pt,labelsep=5pt, topsep=0pt,parsep=0pt,partopsep=0pt,noitemsep

### v1

@@ A.2 GLUE scores of GPT-2
Table 9: Comparison between GLUE scores of different pre-train methods on GPT-2 models. We use full parameters to evaluate dense and SR-STE+DF models, and half parameters (2:4 sparse) to evaluate S-STE models.
params method val loss avg score cola mnli mrpc qnli qqp rte sst2 stsb wnli
124M Dense 2.907 $73.9\pm 1.1$ $44.6\pm 0.9$ $82\pm 0.1$ $78.3\pm 1.3/84.8\pm 1$ $88.4\pm 0.2$ $90\pm 0$ $86.5\pm 0/61.3\pm 1.5$ $91.9\pm 0.2$ $77.3\pm 3.2/77.9\pm 2.9$ $24.3\pm 7.1$
SR-STE $3.0763$ $72.6\pm 0.2$ $41.9\pm 0.3$ $81.0\pm 0.2$ $76.3\pm 0.9/83.4\pm 0.7$ $87.0\pm 0.3$ $89.3\pm 0.1$ $85.6\pm 0.1/60.6\pm 3.4$ $90.9\pm 0.4$ $76.2\pm 3.2/76.5\pm 3.0$ $21.8\pm 4.4$
S-STE $2.984$ $74.1\pm 0.4$ $42.3\pm 1.1$ $80.5\pm 2.8$ $79.3\pm 1.9/85.6\pm 1.4$ $88.1\pm 0.2$ $89.8\pm 0.1$ $86.2\pm 0.1/62.9\pm 1.1$ $91.9\pm 0.4$ $81.0\pm 1.1/81.2\pm 1$ $20.8\pm 3.6$
350M Dense 2.618 $76.3\pm 0.1$ $54.3\pm 0.4$ $85.1\pm 0.1$ $80.7\pm 1/86.6\pm 0.7$ $90.7\pm 0.1$ $91\pm 0.1$ $87.8\pm 0.1/64.9\pm 1.7$ $93.5\pm 0.4$ $81.7\pm 1.2/82.2\pm 0.8$ $17.6\pm 3.2$
SR-STE $2.7180$ $76.3\pm 0.4$ $50.0\pm 1.7$ $84.1\pm 0.2$ $81.4\pm 1.5/87.1\pm 1.1$ $90.0\pm 0.3$ $90.6\pm 0.1$ $87.3\pm 0.1/67.9\pm 1.5$ $93.3\pm 0.4$ $81.3\pm 1.5/81.4\pm 1.4$ $20.6\pm 3.8$
S-STE $2.713$ $76.9\pm 0.6$ $54.2\pm 1.7$ $84.6\pm 0.2$ $80.2\pm 1.3/86.1\pm 0.9$ $90.5\pm 0.3$ $90.8\pm 0.1$ $87.5\pm 0.2/65.1\pm 1.9$ $93.7\pm 0.4$ $83.6\pm 1.1/83.8\pm 1.1$ $22.5\pm 3.9$
774M Dense 2.493 $76.2\pm 0.4$ $57.5\pm 2$ $86.1\pm 0.1$ $80.3\pm 1.3$/$86.4\pm 0.9$ $91.4\pm 0.2$ $91.1\pm 0.1$ $88\pm 0.1$/$67.7\pm 2.6$ $94.6\pm 0.4$ $77.3\pm 3.3$/$78.4\pm 2.9$ $15.1\pm 2.3$
SR-STE+DF $2.564$ $77.1\pm 0.4$ $55.9\pm 0.9$ $85.6\pm 0.2$ $81.2\pm 0.6/87\pm 0.4$ $91.4\pm 0.1$ $91\pm 0.1$ $87.8\pm 0.1/71.5\pm 0.7$ $94.2\pm 0.4$ $81.8\pm 1.3/82.3\pm 1.2$ $15.8\pm 1.2$
S-STE $2.547$ $78.6\pm 0.8$ $57.3\pm 2.7$ $86.6\pm 0.2$ $80.6\pm 1.4/86.6\pm 0.9$ $92\pm 0.1$ $91.5\pm 0.1$ $88.5\pm 0.1/78.3\pm 1.5$ $94.9\pm 0.3$ $85.5\pm 1.2/85.7\pm 1.1$ $16.1\pm 5.9$
A.3 Limitations
@@
A.4 Broader Impact
A.5 Experiments compute resources
See Table 11.

### v3

@@ 8 Conclusions and future work
We would like to thank Ziteng Wang, Bingrui Li, Haocheng Xi, Kang Zhao and Jintao Zhang, Brian Chmiel and Daniel Soudry for valuable discussions. This work was supported by the National Key Research and Development Program of China (No. 2021ZD0110502) and NSFC Project (Nos. 62376131). J.Z is also supported by the XPlorer Prize.
@@ A.2 GLUE scores of GPT-2
Table 9: Comparison between GLUE scores of different pre-train methods on GPT-2 models. This table is the elaboration of Table 6.
Params Method Avg score CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B WNLI
124M Dense $73.9\pm 1.1$ $44.6\pm 0.9$ $82.0\pm 0.1$ $78.3\pm 1.3/84.8\pm 1.0$ $88.4\pm 0.2$ $90.0\pm 0.0$ $86.5\pm 0.0/61.3\pm 1.5$ $91.9\pm 0.2$ $77.3\pm 3.2/77.9\pm 2.9$ $24.3\pm 7.1$
T-SR-STE+DF Hu et al. (2024) $74.3\pm 0.5$ $44.8\pm 1.3$ $81.5\pm 0.2$ $77.5\pm 1.8/84.2\pm 1.3$ $87.8\pm 0.1$ $89.5\pm 0.1$ $85.9\pm 0.1/66.0\pm 1.0$ $90.6\pm 0.4$ $80.0\pm 0.8/80.3\pm 0.5$ $23.9\pm 6.4$
T-SR-STE $72.6\pm 0.2$ $41.9\pm 0.3$ $81.0\pm 0.2$ $76.3\pm 0.9/83.4\pm 0.7$ $87.0\pm 0.3$ $89.3\pm 0.1$ $85.6\pm 0.1/60.6\pm 3.4$ $90.9\pm 0.4$ $76.2\pm 3.2/76.5\pm 3.0$ $21.8\pm 4.4$
SR-STE $73.8\pm 0.3$ $38.3\pm 2.8$ $80.9\pm 0.2$ $79.7\pm 0.7/85.9\pm 0.6$ $87.1\pm 0.3$ $89.5\pm 0.1$ $85.8\pm 0.2/65.5\pm 1.0$ $90.5\pm 0.4$ $80.9\pm 1.2/80.9\pm 1.1$ $20.1\pm 1.8$
S-STE $74.1\pm 0.4$ $42.3\pm 1.1$ $80.5\pm 2.8$ $79.3\pm 1.9/85.6\pm 1.4$ $88.1\pm 0.2$ $89.8\pm 0.1$ $86.2\pm 0.1/62.9\pm 1.1$ $91.9\pm 0.4$ $81.0\pm 1.1/81.2\pm 1.0$ $20.8\pm 3.6$
350M Dense $76.3\pm 0.1$ $54.3\pm 0.4$ $85.1\pm 0.1$ $80.7\pm 1.0/86.6\pm 0.7$ $90.7\pm 0.1$ $91.0\pm 0.1$ $87.8\pm 0.1/64.9\pm 1.7$ $93.5\pm 0.4$ $81.7\pm 1.2/82.2\pm 0.8$ $17.6\pm 3.2$
T-SR-STE+DF Hu et al. (2024) $77.1\pm 0.2$ $51.8\pm 1.8$ $84.3\pm 0.1$ $80.6\pm 1.3/86.5\pm 0.8$ $90.4\pm 0.2$ $90.7\pm 0.1$ $87.5\pm 0.1/66.7\pm 1.3$ $93.3\pm 0.4$ $83.4\pm 1.1/83.5\pm 1.1$ $26.4\pm 4.0$
T-SR-STE $76.3\pm 0.4$ $50.0\pm 1.7$ $84.1\pm 0.2$ $81.4\pm 1.5/87.1\pm 1.1$ $90.0\pm 0.3$ $90.6\pm 0.1$ $87.3\pm 0.1/67.9\pm 1.5$ $93.3\pm 0.4$ $81.3\pm 1.5/81.4\pm 1.4$ $20.6\pm 3.8$
SR-STE $76.8\pm 0.4$ $47.2\pm 3.0$ $84.3\pm 0.2$ $81.4\pm 0.9/87.2\pm 0.6$ $90.2\pm 0.1$ $90.8\pm 0.1$ $87.6\pm 0.1/68.3\pm 1.4$ $93.9\pm 0.1$ $82.0\pm 1.6/82.0\pm 1.7$ $27.1\pm 3.1$
S-STE $76.9\pm 0.6$ $54.2\pm 1.7$ $84.6\pm 0.2$ $80.2\pm 1.3/86.1\pm 0.9$ $90.5\pm 0.3$ $90.8\pm 0.1$ $87.5\pm 0.2/65.1\pm 1.9$ $93.7\pm 0.4$ $83.6\pm 1.1/83.8\pm 1.1$ $22.5\pm 3.9$
774M Dense $76.2\pm 0.4$ $57.5\pm 2.0$ $86.1\pm 0.1$ $80.3\pm 1.3/86.4\pm 0.9$ $91.4\pm 0.2$ $91.1\pm 0.1$ $88.0\pm 0.1/67.7\pm 2.6$ $94.6\pm 0.4$ $77.3\pm 3.3/78.4\pm 2.9$ $15.1\pm 2.3$
T-SR-STE+DF Hu et al. (2024) $77.1\pm 0.4$ $55.9\pm 0.9$ $85.6\pm 0.2$ $81.2\pm 0.6/87.0\pm 0.4$ $91.4\pm 0.1$ $91.0\pm 0.1$ $87.8\pm 0.1/71.5\pm 0.7$ $94.2\pm 0.4$ $81.8\pm 1.3/82.3\pm 1.2$ $15.8\pm 1.2$
S-STE $78.6\pm 0.8$ $57.3\pm 2.7$ $86.6\pm 0.2$ $80.6\pm 1.4/86.6\pm 0.9$ $92.0\pm 0.1$ $91.5\pm 0.1$ $88.5\pm 0.1/78.3\pm 1.5$ $94.9\pm 0.3$ $85.5\pm 1.2/85.7\pm 1.1$ $16.1\pm 5.9$
A.3 Acceleration
See Table 10.
Table 10: Pre-training acceleration ratio with different different batch size $N$, sequence length $n$, embedding dimension $d$ and heads number $h$ on single FFN block and transformer block of GPT-2 with RTX 3090 GPUs.
N n d h FFN GPT-2
Pre-train 4 2048 5120 40 1.31 1.18
16 2048 7168 56 1.32 1.18
8 2048 7168 56 1.33 1.17
4 2048 7168 56 1.31 1.17
4 2048 9216 72 1.31 1.18
Inference 16 2048 7168 56 1.54 1.23
8 2048 7168 56 1.46 1.15
A.4 Limitations
@@
A.5 Broader Impact
A.6 Experiments compute resources
To replicate our experiments, we provide the estimated GPU hours of each setting; see Table 12.